OpenStudy (anonymous):

Find the indefinite integral of: (2x^(3/2)+1)x^(1/2) I keep getting (1/6)(2x^(3/2)+1)^2 + C, but this is incorrect.

OpenStudy (amistre64):

that is definantly an eyesore

OpenStudy (amistre64):

multiply it thru to get seperate terms for addition right?

OpenStudy (amistre64):

2x^(4/2) + x^(1/2)

OpenStudy (amistre64):

[S] 2x^2 + x^(1/2) dx 2x^3/3 + 2x^(3/2)/3 + C right?

OpenStudy (amistre64):

\[\frac{2x^3}{3} +\frac{2x^ \frac{3}{2}}{3} + C\]

OpenStudy (anonymous):

Yes, I believe that is right. I was trying to over-complicate things by using substitution somehow. Thank you.

OpenStudy (amistre64):

youre welcome :) simplify when needed right :)